INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF PLATINUM SULFIDE IN THE NEAR IN-FRARED

\underline{JAMES J. O’BRIEN}, Department of Chemistry and Biochemistry and Center for Nanoscience, University of Missouri, St Louis, MO 63121-4499; LEAH C. O’BRIEN, KIMBERLY HANDLER, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

A new electronic transition of PtS has been recorded using intracavity laser absorption spectroscopy. Two branches associated with a red-degraded bandhead at 12460cm^{−1} are identified as the (0,0) band of a new [12.5] ^{1}Σ^{+} - X ^{1}Σ^{+} transition. The results of the analysis will be presented.